General Meeting Summary 2/15/17

From cua_phy
Jump to navigation Jump to search

JLAB DETECTORS - KPP RUNS (Arthur)

  • SHMS aerogel detector is ready for KPP runs (see: Picture 1 and Picture 2 for signals)
  • Source of attentuation was a splitter (signal split: 50% into fADC and 50% into trigger) upstairs in the counting house that was added for including aerogel detector in the trigger - this will not be the case, so removed the splitter, now signal is fine
  • Problem with first channel not counting was a loose base on the PMT - perhaps happened when people worked on other detectors, after pushing base back in place channel started counting
  • Notice that there is still some noise in the spectra - not clear where it comes from, seen in all detectors. However, aerogel signal is big enough that this level of noise is tolerable
  • Summing of signals - software only, will be addressed later
  • SP-11 tray: Arthur checking with Walter


DIRC DETECTOR WORK (Greg)

  • Work on monitoring system for moving Babar DIRC boxes to JLab is ongoing - found in initial tests that each bar will have to be monitored separately
  • NLAK (material used for lens in EIC DIRC focusing system) radiation resistance tests were performed with help from Marco and Lee Allison (ODU)
  • result shows that material is not radiation hard and need to find another one
  • discussion about using LYSO/LSO, which was also considered for calorimetry, but rejected due to cost


PIXEL DETECTORS (Rachel)

  • Barrel pixel detector has been tested
  • optimizations ongoing for new power up sequence
  • Discussion about Rachel's recent presentation to promote exchange with the semiconductor industry
  • one aspect discussed was radiation testing of components at CERN
  • Discussion about CUA pixel detector lab readiness


PWO CRYSTAL LIGHT YIELD TEMPERATURE DEPENDENCE (Salina)

  • Measured light yields between -10 degC and +18degC - assumption is that crystal temperature is equal to freezer temperature
  • Tested a method of measuring crystal temperature with a laser thermometer - at -10degC the results were unreliable. Next, will try measurements at a temperature closer to room temperature


PWO CRYSTAL PROTOTYPE (Richard)

  • Design ongoing - stepper motor assembled


TCS UPDATE

  • Discussion about a presentation at Thursday's dilepton meeting - Vardan will do it. Slides should include a summary of the two photon source options and TCS measurement with such a photon source
  • Discussion about comparison of Hall C projection with HDICE/Hall B
  • sources of uncertainties: photon flux for Hall C
  • note that HDICE includes assumptions on: luminosity, polarization target and duty factor


NEXT MEETING: WED. 22 FEBRUARY at 10:00 AM (EST)